RITZ ASSIGNMENT AND EVALUATION OF THE WATSONIAN PARAMETERS FOR THE ν_{1} N−H STRETCHING BAND OF DIFLUOROMETHANIMINE, F_{2}C=NH

ALBERTO GAMBI, Dipartimento di Scienze e Tecnologie Chimiche dell’Università di Udine, Via Cotonificio 108, I-33100 Udine, Italy; \underline{GIOVANNI MORUZZI}, Dipartimento di Fisica “Enrico Fermi” dell’Università di Pisa and INFM, Via Filippo Buonarroti 2, I-56127 Pisa, Italy; GOTTFRIED PAWELKE AND HANS BÜRGER, Anorganische Chemie, FB 9, Universität-Gesamthochschule, D-42097 Wuppertal, Germany.

Difluoromethanimine, F_{2}C=NH, is a planar, asymmetric rotor close to the oblate symmetric top limit, with κ = 0.8928. The first successful synthesis of this molecule was carried out by Bürger and Pawelke^{a} in 1988. In the present work we present an investigation of the in-plane ν_{1} fundamental, N−H stretching, centered at 3042.3 cm^{−1}; this band, of symmetry species A’, presents a hybrid ab-type contour with a prominent Q head and wide P and R branches. The difficulties encountered in the assignment of this band are due both to systematic line overlappings and to interactions with dark states — consisting of combination bands in which at least three vibrational quanta are involved — these interactions being particularly strong for 13 ≤ K_{c} ≤ 16. The capabilities of the “Multi-Molecule Ritz Program”, presented at this conference in 1996,^{b} have been enhanced by adding a “feed-back loop” which calculates the best-fit Watsonian parameters directly from the Ritz data-base. This has led to a list of approximately 2300 assigned lines belonging to the investigated band. These assignments, and the perturbations encountered, will be discussed.

aH. Bürger and G. Pawelke, J. Chem. Soc., Chem. Commun., 1988, 105

bG. Moruzzi, Ohio State University 51st International Symposium on Molecular Spectroscopy, TC06, p. 119 (1996).